OpenStudy (anonymous):

Find the value of 2 times the square root of 3, all over the square root of 10. in simplest form.

OpenStudy (jhannybean):

\[\frac{ 2\sqrt{3} }{ \sqrt{10} }\] So when you have a radical in denominator, you have to multiply both top and bottom by the radical

OpenStudy (jhannybean):

You get \[\frac{ (2\sqrt{3}*\sqrt{10} )}{ \sqrt{10} *\sqrt{10} }\]

OpenStudy (jhannybean):

Are you able to reduce it from there?

OpenStudy (jhannybean):

\[\frac{ 2\sqrt{30} }{ 10}=\frac{ \sqrt{30} }{ 5 }\]